Recent information published by the U.S. Environmental Protection Agency indicates that Honda is the manufacturer of four of the top nine vehicles in terms of fuel economy.
a. Determine the probability distribution for the number of Hondas in a sample of three cars chosen from the top nine.

b. What is the likelihood that in the sample of three at least one Honda is included?

Answer by stanbon(75887)

Recent information published by the U.S. Environmental Protection Agency indicates that Honda is the manufacturer of four of the top nine vehicles in terms of fuel economy.
-----
Binomial condition with p = 4/9
-----------
a. Determine the probability distribution for the number of Hondas in a sample of three cars chosen from the top nine.
P(x = 0 hondas) = (5/9)^3
P(x = 1 honda) = 3C1(4/9)(5/9)^2
P(x = 2 hondas) = 3C2(4/9)^2(5/9)
P(x = 3 hondas) = (4/9)^3
===================================
b. What is the likelihood that in the sample of three at least one Honda is included?
P(at least one) = 1 - P(none)
= 1 - (5/9)^3
====================
Cheers,
Stan H.
====================